# Rate doubly robust estimation for weighted average treatment effects

### v1

@@ A.1 Preliminaries and Lemmata
$\quad-\mathbb{P}\left[\frac{\lambda\{\hat{e}(\mathbf{X})\}}{1-\hat{e}(\mathbf{X})}\{\hat{e}(\mathbf{X})-e(\mathbf{X})\}\{\mu_{0}(\mathbf{X})-\hat{\mu}_{0}(\mathbf{X})\}\right]+\mathbb{P}[\tau(\mathbf{X})\{\lambda\{\hat{e}(\mathbf{X})\}-\lambda\{e(\mathbf{X})\}-\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{e}(\mathbf{X})-e(\mathbf{X})\}\}]$
$\quad+\mathbb{P}[\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{\tau}(\mathbf{X})-\tau(\mathbf{X})\}\{e(\mathbf{X})-\hat{e}(\mathbf{X})\}],$
@@
$\quad+\|\mathbb{P}[\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{\tau}(\mathbf{X})-\tau(\mathbf{X})\}\{e(\mathbf{X})-\hat{e}(\mathbf{X})\}]\|_{2}+\|\mathbb{P}[\tau(\mathbf{X})\{\lambda\{\hat{e}(\mathbf{X})\}-\lambda\{e(\mathbf{X})\}-\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{e}(\mathbf{X})-e(\mathbf{X})\}\}]\|_{2}.$ (A.3.9)
@@ B.1.1 Theory of Gateaux derivatives
$\quad+\dot{\lambda}\{e_{r}(\mathbf{X})\}\left[\frac{\partial}{\partial r}\{\mu_{1,r}(\mathbf{X})-\mu_{0,r}(\mathbf{X})\}\right]\{A-e_{r}(\mathbf{X})\}+\dot{\lambda}\{e_{r}(\mathbf{X})\}\{\mu_{1,r}(\mathbf{X})-\mu_{0,r}(\mathbf{X})\}\frac{\partial}{\partial r}\{A-e_{r}(\mathbf{X})\},$ (B.1.4)
@@
$\frac{\partial}{\partial r}\psi_{\tau}(\mathbf{V};\phi_{r})$ $=\{Y-\mu_{1,r}(\mathbf{X})\}\cdot\frac{\partial}{\partial r}\left\{\frac{A}{e_{r}(\mathbf{X})}\right\}-\frac{A}{e_{r}(\mathbf{X})}\cdot\frac{\partial}{\partial r}\mu_{1,r}(\mathbf{X})$
$\quad-\{Y-\mu_{0,r}(\mathbf{X})\}\cdot\frac{\partial}{\partial r}\left\{\frac{1-A}{1-e_{r}(\mathbf{X})}\right\}$
@@
$\quad+\frac{1-A}{1-e_{r}(\mathbf{X})}\{\tilde{\mu}_{0}(\mathbf{X})-\mu_{0}(\mathbf{X})\}+[\{\tilde{\mu}_{1}(\mathbf{X})-\mu_{1}(\mathbf{X})\}-\{\tilde{\mu}_{0}(\mathbf{X})-\mu_{0}(\mathbf{X})\}],$ (B.1.6)
therefore
@@
$=\mathbb{E}[\dot{\lambda}\{e_{r}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}\{\mu_{1,r}(\mathbf{X})-\mu_{0,r}(\mathbf{X})\}]-r\cdot\mathbb{E}\left[\dot{\lambda}\{e_{r}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}\frac{e(\mathbf{X})}{e_{r}(\mathbf{X})}\{\tilde{\mu}_{1}(\mathbf{X})-\mu_{1}(\mathbf{X})\}\right]$
@@
B.1.2 Assumptions and two important lemmata in (2018)
In this section, we introduce two lemmata by [29], which summarized the conditions for asymptotic normality of DML estimators. First, the following assumptions are needed, following the notations of Gateaux derivatives in Section B.1.1. Suppose $C_{1}\geq C_{0}>0$ are some finite constants, $\{\delta_{n}\}_{n\geq 1}$ and $\{\Delta_{n}\}_{n\geq 1}$ are some sequences of positive sequence converging to 0 with $\delta_{n}\geq n^{-1/2}$.
Assumption B.1.3 (Assumption 3.1 in [29]).
@@ B.1.2 Assumptions and two important lemmata in (2018)
$\leq\left(\mathbb{E}\left|\lambda\{e(\mathbf{X})\}\right|^{q}\right)^{1/q}+\left(\mathbb{E}\left[\left|e(\mathbf{X})-\tilde{e}(\mathbf{X})\right|^{q}\cdot\left|\dot{\lambda}\{\bar{e}(\mathbf{X})\}-\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\right|^{q}\right]\right)^{1/q}+2\sup_{\mathbf{X}}\left|\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\right|$
@@
$\quad+\left(\mathbb{E}\big|\{\tilde{\mu}_{1}(\mathbf{X})-\tilde{\mu}_{0}(\mathbf{X})\}\cdot e(\mathbf{X})\cdot[\dot{\lambda}\{\bar{e}(\mathbf{X})\}-\dot{\lambda}\{\tilde{e}(\mathbf{X})\}]\big|^{q}\right)^{1/q}+\left(\mathbb{E}\big|\{\tilde{\mu}_{1}(\mathbf{X})-\tilde{\mu}_{0}(\mathbf{X})\}\cdot\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\cdot\{A-e(\mathbf{X})\}\big|^{q}\right)^{1/q}$
@@
$=\|\dot{\lambda}\{\bar{e}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}-A\ddot{\lambda}\{\bar{e}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}+\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}+e(\mathbf{X})\{\dot{\lambda}\{\tilde{e}(\mathbf{X})\}-\dot{\lambda}\{e(\mathbf{X})\}\}\|_{2}$
@@
$\mathcal{I}_{17}$ $\leq\sup_{\mathbf{X}}\left|\frac{{\lambda\{\tilde{e}(\mathbf{X})\}}/\{1-\tilde{e}(\mathbf{X})\}-{\lambda\{e(\mathbf{X})\}}/\{1-e(\mathbf{X})\}}{\tilde{e}(\mathbf{X})-e(\mathbf{X})}\right|\cdot\sup_{\mathbf{X}}\left(\mathbb{E}[(1-A)\{Y-\mu_{0}(\mathbf{X})\}^{2}\mid\mathbf{X}]\right)^{1/2}\cdot\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}$
@@
$\quad+\|\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\cdot\tilde{e}(\mathbf{X})\cdot[\{\tilde{\mu}_{1}(\mathbf{X})-\tilde{\mu}_{0}(\mathbf{X})\}-\{\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})\}]\|_{2}+\|[\dot{\lambda}\{e(\mathbf{X})\}e(\mathbf{X})-\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\tilde{e}(\mathbf{X})]\cdot\tau(\mathbf{X})\|_{2}$
@@
Furthermore, by Conditions 11 and 10 as well as plugging (B.2.5) and (B.2.11) into B.2.4, we get an upper bound on $r_{n}^{\prime}$ as $r_{n}^{\prime}\preceq\left|\gamma_{0}\right|\cdot\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}+\sum_{a=0}^{1}\|\tilde{\mu}_{a}(\mathbf{X})-\mu_{a}(\mathbf{X})\|_{2}.$ Finally, we bound $\lambda^{\prime}_{n}$ in (c) in Assumption B.1.4 by Lemma B.1.2 and Cauchy-Schwarz inequality:
@@
$\preceq\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}\cdot\sum_{a=0}^{1}\|\mu_{a}(\mathbf{X})-\mu_{a}(\mathbf{X})\|_{2}+\sup_{r\in[0,1]}\left\{\|\lambda^{(3)}\{e_{r}(\mathbf{X})\}\|_{2}\cdot\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}+\|\ddot{\lambda}\{e_{r}(\mathbf{X})\}\|_{2}\right\}\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}^{2},$

### v2

@@ A.1 Preliminaries and Lemmata
$\quad-\mathbb{P}\left[\frac{\lambda\{\hat{e}(\mathbf{X})\}}{1-\hat{e}(\mathbf{X})}\{\hat{e}(\mathbf{X})-e(\mathbf{X})\}\{\mu_{0}(\mathbf{X})-\hat{\mu}_{0}(\mathbf{X})\}\right]+\mathbb{P}[\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{\tau}(\mathbf{X})-\tau(\mathbf{X})\}\{e(\mathbf{X})-\hat{e}(\mathbf{X})\}]$
$\quad+\mathbb{P}[\tau(\mathbf{X})\{\lambda\{\hat{e}(\mathbf{X})\}-\lambda\{e(\mathbf{X})\}-\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{e}(\mathbf{X})-e(\mathbf{X})\}\}],$
@@
$\quad+\|\mathbb{P}[\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{\tau}(\mathbf{X})-\tau(\mathbf{X})\}\{e(\mathbf{X})-\hat{e}(\mathbf{X})\}]\|_{2}$
$\quad+\|\mathbb{P}[\tau(\mathbf{X})\{\lambda\{\hat{e}(\mathbf{X})\}-\lambda\{e(\mathbf{X})\}-\dot{\lambda}\{\hat{e}(\mathbf{X})\}\{\hat{e}(\mathbf{X})-e(\mathbf{X})\}\}]\|_{2}.$ (A.3.9)
@@ B.1.1 Theory of Gateaux derivatives
$\quad+\dot{\lambda}\{e_{r}(\mathbf{X})\}\left[\frac{\partial}{\partial r}\{\mu_{1,r}(\mathbf{X})-\mu_{0,r}(\mathbf{X})\}\right]\{A-e_{r}(\mathbf{X})\}$
$\quad+\dot{\lambda}\{e_{r}(\mathbf{X})\}\{\mu_{1,r}(\mathbf{X})-\mu_{0,r}(\mathbf{X})\}\frac{\partial}{\partial r}\{A-e_{r}(\mathbf{X})\},$ (B.1.4)
@@
$\frac{\partial}{\partial r}\psi_{\tau}(\mathbf{V};\phi_{r})$ $=\{Y-\mu_{1,r}(\mathbf{X})\}\cdot\frac{\partial}{\partial r}\left\{\frac{A}{e_{r}(\mathbf{X})}\right\}-\frac{A}{e_{r}(\mathbf{X})}\cdot\frac{\partial}{\partial r}\mu_{1,r}(\mathbf{X})-\{Y-\mu_{0,r}(\mathbf{X})\}\cdot\frac{\partial}{\partial r}\left\{\frac{1-A}{1-e_{r}(\mathbf{X})}\right\}$
@@
$\quad+\frac{1-A}{1-e_{r}(\mathbf{X})}\{\tilde{\mu}_{0}(\mathbf{X})-\mu_{0}(\mathbf{X})\}+[\{\tilde{\mu}_{1}(\mathbf{X})-\mu_{1}(\mathbf{X})\}-\{\tilde{\mu}_{0}(\mathbf{X})-\mu_{0}(\mathbf{X})\}].$ (B.1.6)
Therefore,
@@
$=\mathbb{E}[\dot{\lambda}\{e_{r}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}\{\mu_{1,r}(\mathbf{X})-\mu_{0,r}(\mathbf{X})\}]$
$\quad-r\cdot\mathbb{E}\left[\dot{\lambda}\{e_{r}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}\frac{e(\mathbf{X})}{e_{r}(\mathbf{X})}\{\tilde{\mu}_{1}(\mathbf{X})-\mu_{1}(\mathbf{X})\}\right]$
@@
B.1.2 Assumptions and two useful lemmata
In this section, we introduce two lemmata by Chernozhukov et al. [29], which summarized the conditions for asymptotic normality of DML estimators. First, the following assumptions are needed, following the notations of Gateaux derivatives in Section B.1.1. Suppose $C_{1}\geq C_{0}>0$ are some finite constants, $\{\delta_{n}\}_{n\geq 1}$ and $\{\Delta_{n}\}_{n\geq 1}$ are some sequences of positive sequence converging to 0 with $\delta_{n}\geq n^{-1/2}$.
Assumption B.1.3 (Assumption 3.1 in Chernozhukov et al. [29]).
@@ B.1.2 Assumptions and two useful lemmata
$\leq\left(\mathbb{E}\left|\lambda\{e(\mathbf{X})\}\right|^{q}\right)^{1/q}+\left(\mathbb{E}\left[\left|e(\mathbf{X})-\tilde{e}(\mathbf{X})\right|^{q}\cdot\left|\dot{\lambda}\{\bar{e}(\mathbf{X})\}-\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\right|^{q}\right]\right)^{1/q}$
$\qquad+2\sup_{\mathbf{X}}\left|\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\right|$
@@
$\quad+\left(\mathbb{E}\big|\{\tilde{\mu}_{1}(\mathbf{X})-\tilde{\mu}_{0}(\mathbf{X})\}\cdot e(\mathbf{X})\cdot[\dot{\lambda}\{\bar{e}(\mathbf{X})\}-\dot{\lambda}\{\tilde{e}(\mathbf{X})\}]\big|^{q}\right)^{1/q}$
$\quad+\left(\mathbb{E}\big|\{\tilde{\mu}_{1}(\mathbf{X})-\tilde{\mu}_{0}(\mathbf{X})\}\cdot\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\cdot\{A-e(\mathbf{X})\}\big|^{q}\right)^{1/q}$
@@
$=\|\dot{\lambda}\{\bar{e}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}-A\ddot{\lambda}\{\bar{e}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}+\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\{\tilde{e}(\mathbf{X})-e(\mathbf{X})\}$
$\quad+e(\mathbf{X})\{\dot{\lambda}\{\tilde{e}(\mathbf{X})\}-\dot{\lambda}\{e(\mathbf{X})\}\}\|_{2}$
@@
$\mathcal{I}_{17}$ $\leq\sup_{\mathbf{X}}\left|\frac{{\lambda\{\tilde{e}(\mathbf{X})\}}/\{1-\tilde{e}(\mathbf{X})\}-{\lambda\{e(\mathbf{X})\}}/\{1-e(\mathbf{X})\}}{\tilde{e}(\mathbf{X})-e(\mathbf{X})}\right|$
$\qquad\times\sup_{\mathbf{X}}\left(\mathbb{E}[(1-A)\{Y-\mu_{0}(\mathbf{X})\}^{2}\mid\mathbf{X}]\right)^{1/2}\cdot\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}$
@@
$\quad+\|\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\cdot\tilde{e}(\mathbf{X})\cdot[\{\tilde{\mu}_{1}(\mathbf{X})-\tilde{\mu}_{0}(\mathbf{X})\}-\{\mu_{1}(\mathbf{X})-\mu_{0}(\mathbf{X})\}]\|_{2}$
$\qquad+\|[\dot{\lambda}\{e(\mathbf{X})\}e(\mathbf{X})-\dot{\lambda}\{\tilde{e}(\mathbf{X})\}\tilde{e}(\mathbf{X})]\cdot\tau(\mathbf{X})\|_{2}$
@@
Furthermore, by Conditions 11 and 10 as well as plugging (B.2.5) and (B.2.11) into B.2.4, we get an upper bound on $r_{n}^{\prime}$ as $r_{n}^{\prime}\preceq\left|\gamma_{0}\right|\cdot\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}+\sum_{a=0}^{1}\|\tilde{\mu}_{a}(\mathbf{X})-\mu_{a}(\mathbf{X})\|_{2}.$
Finally, we bound $\lambda^{\prime}_{n}$ Assumption B.1.4(c) by Lemma B.1.2 and Cauchy-Schwarz inequality:
@@
$\preceq\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}\cdot\sum_{a=0}^{1}\|\mu_{a}(\mathbf{X})-\mu_{a}(\mathbf{X})\|_{2}$
$\quad+\sup_{r\in[0,1]}\left\{\|\lambda^{(3)}\{e_{r}(\mathbf{X})\}\|_{2}\cdot\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}+\|\ddot{\lambda}\{e_{r}(\mathbf{X})\}\|_{2}\right\}\|\tilde{e}(\mathbf{X})-e(\mathbf{X})\|_{2}^{2},$